OpenStudy (anonymous):

okay so Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1000 times the wavelength of an X-ray. find the wavelength of ultraviolet rays? okay! soo the X-ray is =10^-10 okay! so I have to find the ultra. so does that mean they just wants want me to fins the scientific notation to 1000? which is 1.0 x 10^3 or do they want me to add or subtract both with each other?

OpenStudy (anonymous):

wavelength of ultraviolet rays=1.0*10^-10 *10^3=1.0*10^(-10+3)=1.0*10^-7